# Effective Fine-Tuning of Vision-Language Models for Accurate Galaxy Morphology Analysis

### v1

@@ 4.2 Feature Projections
Figure 6: Performance Comparison of Different Methods on MNIST and Galaxy10 Datasets (F1 Score)
4.3 Galaxy Morphology Classification
@@
The self-supervised models (MAE [10], DINOv2 [19], and MSN [1]), when fine-tuned following Lastufka et al. [27], show promising performance on GalaxyMNIST but are generally outperformed by models pretrained with supervised learning [8, 9], particularly on Galaxy10. This result suggests that while self-supervised pretraining captures general features, supervised models pretrained on natural images, tend to transfer more effectively to the morphological distinctions in astronomical images.
The Zoobot models [32] (using MaxViT and ConvNeXT as the backbone network, respectively), specifically pretrained on large-scale galaxy data, demonstrate strong performance on both datasets. This result underscores the value of domain-specific pretraining for high-performance astronomical classification.
@@
Figure 7: Examples of similarity search on the Galaxy10 dataset. The top row shows the query galaxy images (outlined in green) for each column, and the other rows display the seven most similar galaxies retrieved by the model.
Figure 6 compares the F1 scores of various models pretrained on natural data across different data sizes for the fine-tuning datasets, GalaxyMNIST and Galaxy10. Our models achieve the highest performance on both datasets, with all data sizes, showing efficient generalization and effective adaptation to astronomical tasks. Self-supervised models (e.g., DINOv2 [19], MAE [10], MSN [1]) and standard supervised models (e.g., ResNet-18, ResNet-50 [9] and ViT [8]) achieve moderate results, indicating that pretraining with natural image alone may not fully capture the domain-specific details required for astronomical tasks. These results illustrate the effectiveness of our multi-modal adaptation strategy in bridging general pretrained models and domain-specific applications in astronomy.
Table 3: Comparison of Mean Average Precision (mAP) on GalaxyMNIST and Galaxy10.
@@
4.4 Similarity Search
In astronomy, automatically identifying the similarity between two galaxies is a challenging but essential task [31]. Effective searches for similar galaxies can help us find counterparts of rare galaxies, making the leap from a one-off discovery to a new class of phenomena [31]. Latent representations encoded by neural networks provide a new opportunity for measuring morphological similarity. In our similarity search evaluation, we measure the morphological similarity of galaxies by comparing feature embeddings through a similarity matrix. Specifically, we calculate pairwise similarities using dot products between feature vectors, creating a matrix where each entry indicates the similarity value between two galaxy images. For each image, we retrieve the most similar images (excluding the image itself) and use these to evaluate retrieval performance.
@@ 4.4 Similarity Search
4.5 Ablation Studies
We conduct an ablation study to evaluate the contribution of each alternative in our approach. First, we compare how our model works with ViT and ConvNeXT. Then we examine the effect of our tri-modality input in comparison with the text-image bimodal CLIP model without the schematic symbol input. Finally, we compare our models across different stages of fine-tuning: Ours_v1 is the warm-up model trained through the first stage, omitting the second training stage; Ours_v2 skips the first stage and directly goes into the second stage of training until convergence; Ours_v3 omits the second stage, training through the first stage to convergence; Ours_Scratch goes through both stages, but is trained from scratch on the Galaxy10 and GalaxyMNIST datasets, without pre-training on large-scale natural datasets.
The results in Table 4 show the impact of each alternative in our approach. The full models, Ours(ViT-16) and Ours(ConvNeXT), achieve the highest accuracy and F1 scores on both datasets. Removing the symbol modality (CLIP models) or omitting training stages (e.g., Ours_v1, Ours_v2, and Ours_v3) reduces performance. The Ours_Scratch variant, trained without pretraining, shows notable drops, indicating the impact of pretraining with large-scale general datasets on astronomy tasks.
Table 4: Ablation Study Results on Classification Task
@@ 4.5 Ablation Studies
5 Conclusion and Future Work
In this paper, we introduced GalaxAlign, a tri-modal framework designed to finetune vision foundation models for galaxy morphology analysis, effectively utilizing schematic symbols, text descriptions, and galaxy images. By integrating domain-specific knowledge in a multi-modal approach, GalaxAlign effectively reuses the general foundation models pre-trained on natural datasets and reduces the need for manually labeled data for training foundation models from scratch.

### v2

@@ 4.2. Feature Projections
Figure 6. Performance Comparison of Different Methods on GalaxyMNIST and Galaxy10 Datasets with Different Data Sizes (F1 Score).
4.3. Galaxy Morphology Classification
@@
The self-supervised models (MAE (He et al., 2022), DINOv2 (Oquab et al., 2023), and MSN (Assran et al., 2022)), when fine-tuned following Lastufka et al. (Voloshynovskyy, 2024), show promising performance on GalaxyMNIST but are generally outperformed by models pretrained with supervised learning (Dosovitskiy, 2020; He et al., 2016), particularly on Galaxy10. This result suggests that finetuning supervised models pretrained on natural images tends to transfer more effectively to the morphological distinctions in astronomical images. And fine-tuning CLIP (Radford et al., 2021) models with images and textual descriptions also achieve better performance than other uni-modal methods.
The Zoobot models (Walmsley et al., 2023a) (using MaxViT and ConvNeXT as backbones, respectively), pretrained on large-scale galaxy data, demonstrate strong performance. This result underscores the value of domain-specific pretraining for astronomical classification.
@@
We also conduct additional evaluation using Gemini-2.5 and QVQ via the latest public APIs. We also finetuned LLaVA1.6 (Liu et al., 2023b) using LoRA (Hu et al., 2022) on our full training datasets. The testing results on these baselines are presented in Appendix.
Table 3. Comparison of Mean Average Precision (mAP) on GalaxyMNIST and Galaxy10.
@@
(a) Few-shot tests on GalaxyMNIST dataset.
(b) Few-shot tests on Galaxy10 dataset.
Figure 7. Few-Shot Tests of Different Methods on GalaxyMNIST and Galaxy10 Datasets.
Figure 8. Examples of similarity search on the Galaxy10 dataset. The top row shows the query galaxy images (outlined in green) for each column, and the other rows display the seven most similar galaxies retrieved by the model.
Figure 6 compares the F1 scores of various models pretrained on natural data across different data sizes for the fine-tuning datasets, GalaxyMNIST and Galaxy10. Our models achieve the highest performance on both datasets, with all data sizes, showing efficient generalization and effective adaptation to astronomical tasks. Self-supervised models (e.g., DINOv2 (Oquab et al., 2023), MAE (He et al., 2022), MSN (Assran et al., 2022)) and standard supervised models (e.g., ResNet-18, ResNet-50 (He et al., 2016) and ViT (Dosovitskiy, 2020)) achieve moderate results, indicating that pretraining with natural image alone may not fully capture the domain-specific details required for astronomical tasks. These results illustrate the effectiveness of our multi-modal adaptation strategy in bridging general pretrained models and domain-specific applications.
4.4. Similarity Search
In astronomy, automatically identifying the similarity between two galaxies is a challenging but essential task (Walmsley et al., 2022b). Effective searches for similar galaxies can help us find counterparts of rare galaxies, making the leap from a one-off discovery to a new class of phenomena (Walmsley et al., 2022b). Latent representations encoded by neural networks provide a new opportunity for measuring morphological similarity. In our similarity search evaluation, we measure the morphological similarity of galaxies by comparing feature embeddings through a similarity matrix. Specifically, we calculate pairwise similarities using dot products between feature vectors, creating a matrix where each entry indicates the similarity value between two galaxy images. For each image, we retrieve the most similar images (excluding the image itself) and use these to evaluate retrieval performance.
@@ 4.4. Similarity Search
4.5. Few-Shot Tests
We conduct few-shot tests on GalaxAlign and other foundation models pretrained on natural data. In Figure 7, across all data sizes, our GalaxAlign model (labeled as ”Ours”) consistently achieves the highest or nearly the highest performance compared to other baseline models, highlighting GalaxAlign’s strong adaptability to the specific characteristics of astronomical images. This advantage is particularly evident on the Galaxy10 dataset, where the increased number of morphological classes better demonstrates the effectiveness of our model.
4.6. Ablation Studies
We conduct an ablation study to evaluate the contribution of each alternative in our approach. First, we compare how well our model works with ViT and ConvNeXT. Then we examine the effect of our tri-modality input in comparison with the text-image bimodal CLIP model without the schematic symbol input. Finally, we compare our models across different stages of fine-tuning: Ours_v1 is the warm-up model trained through the first stage, omitting the second training stage; Ours_v2 skips the first stage and directly goes into the second stage of training until convergence; Ours_v3 omits the second stage, training through the first stage to convergence; Ours_Scratch goes through both stages, but is trained from scratch on the Galaxy10 and GalaxyMNIST datasets, without pre-training on large-scale natural datasets. The results in Table 4 show the impact of each alternative in our approach. The full models, Ours(ViT-16) and Ours(ConvNeXT), achieve the highest accuracy and F1 scores on both datasets. Removing the symbol modality (CLIP models) or omitting training stages (e.g., Ours_v1, Ours_v2, and Ours_v3) reduces performance. The Ours_Scratch variant, trained without pretraining, shows notable drops, indicating the impact of pretraining with large-scale general datasets on astronomy tasks.
We also conduct the ablation study of text and symbols and the ablation study on each part of our tri-model loss function. Results in Table 5 show that removing either the schematic symbols, text descriptions in our framework or each modal pair in the loss function leads to a performance drop. Each modality contributes to our tasks in a complementary way.
Table 4. Ablation Study of Training Porcess
@@ 4.6. Ablation Studies
Table 5. Ablation Study of Modalities and Loss
Model Variant Galaxy10 GalaxyMNIST
Accuracy F1 Accuracy F1
ViT
w/o Symbol 0.9635 0.9549 0.9125 0.9129
w/o Text 0.7965 0.7627 0.9050 0.9053
Loss w/o text-image 0.9617 0.9582 0.9100 0.9102
Loss w/o symbol-image 0.9622 0.9576 0.9125 0.9127
Loss w/o text-symbol 0.9634 0.9587 0.8945 0.9108
Ours (Full) 0.9732 0.9702 0.9272 0.9276
ConvNeXT
w/o Symbol 0.9625 0.9537 0.9310 0.9316
w/o Text 0.7599 0.7330 0.8975 0.8952
Loss w/o text-image 0.5781 0.5503 0.7390 0.7379
Loss w/o symbol-image 0.9552 0.9444 0.9360 0.9364
Loss w/o text-symbol 0.9560 0.9491 0.9340 0.9346
Ours (Full) 0.9710 0.9664 0.9372 0.9377
5. Conclusion and Future Work
In this paper, we introduced GalaxAlign, a tri-modal framework designed to finetune vision foundation models for galaxy morphology analysis, utilizing schematic symbols, text descriptions, and galaxy images. By integrating domain-specific knowledge in a multi-modal approach, GalaxAlign effectively reuses the general foundation models pre-trained on natural datasets and reduces the need for manually labeled data to train foundation models from scratch.
@@ 5. Conclusion and Future Work
Acknowledgements
This work was supported by the China National Science Foundation (NSF) No. 62372393 and a startup fund from HKUST (Guangzhou).
Appendix A Additional Experiment Details
A.1. Setup Details
The ConvNeXT version of our model is trained with Adam with $lr=5e-6,\text{weight decay}=0.0002$. The ViT version of our model is trained with Adam with $lr=1e-4,\text{weight decay}=0.02$. We adopt a full fine-tuning strategy for our method. For the first stage, we train 50 epochs for the best results. For each dataset, we randomly split 20% of all images for testing, with the remainder being used for training. We repeat all experiments five times and compute the mean and standard deviation of the results.
A.2. Baseline Details
To ensure a fair comparison and achieve the best performance for each baseline, we followed the fine-tuning strategies recommended in prior work. Specifically, for MAE (He et al., 2022), DINOv2 (Oquab et al., 2023), MSN (Assran et al., 2022), and ViT-16 (Dosovitskiy, 2020), we adopted the approach outlined by Lastufka et al. (Voloshynovskyy, 2024), which involves freezing the encoder parameters and fine-tuning a classification head. In contrast, for ResNet18, ResNet50 (He et al., 2016), and Zoobot (Walmsley et al., 2023a), we used the same full fine-tuning strategy as our method, where all parameters in the model are updated during training, to maximize their performance on the tasks. This ensures that each baseline is optimized using its most suitable training procedure.
Appendix B Additional Results
B.1. Similarity Search
Figure 9 presents the results of a similarity search performed by GalaxAlign (ConvNeXT) on the GalaxyMNIST dataset. Each column represents a query galaxy (top row, outlined in green), followed by the seven most similar galaxies retrieved by the model (subsequent rows). The results demonstrate the model’s ability to effectively retrieve galaxies with similar visual and morphological features. Across various galaxy types, the top-ranked matches closely resemble the query galaxy’s shape and structure, indicating strong model performance in identifying morphological similarities.
Figure 10. Visualization of the representations learned by our method on the Galaxy10 dataset, illustrating similar galaxies occupying nearby regions in the feature space. This visualization is created using PCA to compress the representation to 2D and place galaxy images at the locations of their corresponding galaxies in the grid.
Figure 11. The t-SNE visualization for features extracted using DINOv2, MAE, MSN, ResNet18, ResNet50, ViT on Galaxy10 dataset.
Figure 12. The t-SNE visualization for features extracted using DINOv2, MAE, MSN, ResNet18, ResNet50, ViT on GalaxyMNIST dataset.
B.2. Additional Comparison with Large Models
We did additional evaluation using Gemini-2.5 and QVQ via the latest public APIs. We also finetuned LLaVA1.6 using LoRA on our full training datasets. The testing results on these baselines are in Table 6. GalaxAlign consistently outperforms these methods, highlighting the importance of our tri-modal approach.
Table 6. Additional Performance Comparison on Galaxy10 and GalaxyMNIST
Model Galaxy10 GalaxyMNIST
Acc F1 Acc F1
Gemini-2.5 0.375 0.239 0.550 0.534
QVQ 0.160 0.170 0.425 0.440
LLaVA1.6 (LoRA) 0.917 0.905 0.896 0.898
B.3. Additional Feature Projections
Figure 13 provides a t-SNE visualization (Van der Maaten and Hinton, 2008) of galaxy data embeddings of GalaxyMNIST dataset learned by different models.
Figure 13. The t-SNE visualization for features extracted using Zoobot (MaxViT and ConvNeXT) and GalaxAlign (ViT-16 and ConvNeXT) GalaxyMNIST dataset.
Figure 11 and Figure 12 illustrate the t-SNE visualizations of features extracted from various baseline models, including DINOv2, MAE, MSN, ResNet18, ResNet50, and ViT16, on the Galaxy10 and GalaxyMNIST datasets, respectively. These visualizations demonstrate the separability of different galaxy classes in the feature space, providing insights into how well the models capture morphological differences among galaxy types.
For the Galaxy10 dataset, ResNet50 and ResNet18 exhibit moderately distinct clusters, indicating their ability to differentiate between galaxy classes. However, the clusters remain partially overlapping, particularly for morphologically similar classes (e.g., Barred Spiral, Unbarred Tight Spiral and Unbarred Loose Spiral). Models such as DINOv2, MAE and MSN show significant class mixing, reflecting suboptimal feature representations. ViT16 demonstrates slightly better clustering compared to DINOv2, MAE and MSN but still struggles to achieve complete separability.
On the GalaxyMNIST dataset, the separability of galaxy classes improves slightly for most models. ResNet18 and ResNet50 again outperform other baselines, showing relatively distinct clusters, although some class overlaps persist. Other models produce dispersed embeddings, with notable mixing between classes, highlighting their limitations in capturing galaxy morphology.
Figure 10 provides a 2D grid visualization of galaxy images sampled from the Galaxy10 dataset. Similar to the GalaxyMNIST visualization, high-dimensional features were reduced using PCA and aligned to a grid. The arrangement highlights the structural similarity between galaxies, with morphologically similar types clustering together in adjacent regions. The smooth transitions between different galaxy classes further demonstrate the ability of our model to effectively capture and encode galaxy morphology in the feature space, enabling clear distinctions and meaningful groupings across diverse galaxy types.